A long distance telephone calling card charges $0.05 per minute, but adds on a $0.75 charge for each call dialed. Using 'C' for the number of calls made and 'I' for the time spent talking, create an expression that will keep track of the cost of using the card.
answer- C=.75 I/.05 ?
Using the expression created calculate the remaining value on a $20.00 phone card after 5 calls that lasted a total of 26 minutes.
answer is $14.95 but don't know how to use an expression to find answer.
I did .75 * 5=3.75 26*.05= 1.30 3.75+1.30=5.05 20.00-5.05= 14.95
Any help would be appreciated as this is my first math course in over 20 years. Answer by nerdybill(7384) 
 A long distance telephone calling card charges $0.05 per minute, but adds on a $0.75 charge for each call dialed. Using 'C' for the number of calls made and 'I' for the time spent talking, create an expression that will keep track of the cost of using the card.
"cost of using card" = "total dialing charges" + "total minutes called"
"cost of using card" = .75C + .05I
.
Using the expression created calculate the remaining value on a $20.00 phone card after 5 calls that lasted a total of 26 minutes.
"remaining cost" = "current value" - "cost"
"remaining cost" = 20 - (.75C + .05I)
"remaining cost" = 20 - (.75(5) + .05(26))
"remaining cost" = 20 - (3.75 + .1.3)
"remaining cost" = 20 - 5.05
"remaining cost" = $14.95